# Fast ECoT: Efficient Embodied Chain-of-Thought via Thoughts Reuse

### v1

@@ Abstract
Embodied Chain-of-Thought (ECoT) reasoning enhances vision-language-action (VLA) models by improving performance and interpretability through intermediate reasoning steps. However, its sequential autoregressive token generation introduces significant inference latency, limiting real-time deployment. We propose Fast ECoT, an inference-time acceleration method that exploits the structured and repetitive nature of ECoT to (1) cache and reuse high-level reasoning across timesteps and (2) parallelise the generation of modular reasoning steps. Additionally, we introduce an asynchronous scheduler that decouples reasoning from action decoding, further boosting responsiveness. Fast ECoT requires no model changes or additional training and integrates easily into existing VLA pipelines. Experiments in both simulation (LIBERO) and real-world robot tasks show up to a 7.5× reduction in latency with comparable or improved task success rate and reasoning faithfulness, bringing ECoT policies closer to practical real-time deployment.
Keywords: Embodied Chain-of-Thought Reasoning, Inference Acceleration, Vision-Language-Action Models
1 Introduction
@@
Building on this reuse, we introduce a partial parallelisation strategy that transforms the traditionally sequential reasoning process into a batched one. By branching only when necessary to handle novel information, Fast ECoT enables multiple reasoning steps to be generated in parallel—substantially reducing inference latency. We further introduce an asynchronous scheduling mechanism that decouples action and reasoning generation. Recognising that robot actions evolve more slowly than sensor observations, we prioritise fast action decoding while allowing reasoning traces to update asynchronously in the background. This design reduces latency without compromising decision quality, as the cached reasoning remains stable over short time horizons.
In summary, our contributions are as follows:
  • We propose Fast ECoT, a novel inference-time acceleration method that caches and reuses intermediate reasoning steps, enabling partially parallel chain-of-thought generation for the first time in embodied policies.
  • We introduce an asynchronous scheduling strategy that decouples reasoning and action generation, allowing high-frequency control signals to be emitted while reasoning is updated in the background.
  • Our method requires no model changes or additional training and can be seamlessly integrated into existing VLA models that perform embodied reasoning, making it broadly usable in practical robotic systems.
2 Related Work
Foundation Models for Robotic Manipulation. Recent advances in robot learning have produced large-scale generalist policies to excel at robotic manipulation tasks, by first pre-training on diverse multimodal datasets [1, 3, 2], and further fine-tuning on extensive robot-specific data collections [5, 6]. Vision-language-action (VLA) architectures [1, 3, 2]—which integrate vision-language models pretrained on internet-scale corpora [7, 8]—further unify perception, language, and control into a single transformer-based policy, demonstrating state-of-the-art performance on benchmarks. Nonetheless, current VLA approaches purely generate robot actions and largely underutilise the inherent multi-step reasoning capabilities of their language backbones.
Reasoning for Robotic Control. Chain-of-thought (CoT) prompting [9, 10] has proven effective in enhancing large language models by encouraging step-by-step reasoning. Prior methods employ pre-trained language models for high-level planning [11, 12, 13], often requiring separate low-level controllers for execution. More recent work explores end-to-end integration of reasoning into robotic policies. Notably, Embodied Chain-of-Thought (ECoT) [4] enables robots to generate grounded, interpretable reasoning traces referencing visual and environmental observations, leading to improved performance when paired with VLA models like OpenVLA [1]. EMMAX [14] similarly embeds reasoning into action pipelines. Our work builds on ECoT, targeting its major bottleneck—high inference latency from autoregressive token generation—by proposing a partially parallelised reasoning framework that preserves interpretability while substantially accelerating inference.
Inference Optimisation in Language and Multimodal Models. A wide range of techniques have been proposed to speed up inference in autoregressive models. Speculative decoding [15, 16] accelerates generation by predicting tokens with a lightweight draft model and verifying them with the full model. Non-autoregressive and parallel decoding strategies have also been adopted, particularly in machine translation and, more recently, in robotic control [17, 18]. Quantisation methods [19, 20, 21] are widely used to reduce model precision for faster computation. In contrast to these methods, which typically operate at the token or model level, our work focuses on reasoning-level acceleration—breaking CoT generation into reusable, semantically meaningful segments that can be cached and executed in parallel. This approach is model-agnostic and does not rely on auxiliary models, fine-tuning, or precision reduction. By enabling efficient reuse and branching during reasoning, our method offers a lightweight and integrable solution for speeding up VLA policies with CoT reasoning, without compromising interpretability or task success.
3 Preliminaries
@@
3.1 Embodied Chain-of-Thought Reasoning (ECoT)
@@
Vision‑Language‑Action (VLA) models [1] build on large pre‑trained vision‑language models, fine‑tuning them to map a natural language instruction $I^{t}$ and image observation $O^{t}$ directly to low‑level robot actions $A^{t}$ via autoregressive token prediction at each control step $t$. Embodied Chain‑of‑Thought (ECoT) [4] reasoning augments this reactive paradigm by supervising the model to generate a structured sequence of $N$ intermediate reasoning steps $R^{t}=\{\,r_{i}^{t}\mid i=1,2,\dots,N\}$, e.g. rephrased task, high‑level plan, grounded sub‑task, low‑level move command, visual features, before emitting the final action $A^{t}$. The reasoning steps are separated into high- and low-level steps, as clarified in Fig. 1 and Sec. A.1.
3.2 Continuous Batching
@@
In autoregressive models, batching improves throughput, but static batching can be inefficient when sequence lengths vary—shorter sequences must wait for longer ones, wasting compute on padding. Continuous batching [22] addresses this by dynamically replacing completed sequences with new ones, maintaining high GPU utilization and minimising idle time. This strategy, used in systems like Tensorrt-LLM [23] and vLLM [24], has shown 2–4× throughput gains in LLM serving.
4 Method
(a) Average update ratio of reasoning steps, highlighting that high‑level reasoning components remain relatively stable, whereas lower‑level steps require more frequent updates.
@@ 4 Method
(b) Average length and standard deviation of different reasoning steps, showing significant disparities in token counts.
@@
Figure 2: Statistics illustrating the pattern of ECoT reasoning steps under Bridge V2 [5].
4.1 Inference Characteristics of ECoT Reasoning
@@
Unlike traditional chain-of-thought approaches in language modelling [25, 26] and vision-language modelling [27, 28], which yield diverse, dynamic reasoning patterns that rarely recur, ECoT consistently follows a structured, periodically recurring workflow: planning, sub-task identification, motion reasoning, and visual-feature processing, before predicting subsequent robot actions. To analyse this behaviour quantitatively, we sample all 1,110 episodes containing reasoning from the Bridge V2 dataset [5] and compute the average and the standard deviation of the update ratio (percentage of reasoning content updated at the next time step) and token length for each reasoning step. As depicted in Fig. 2(a), higher-level reasoning components in ECoT (e.g., planning and subtask reasoning) remain relatively similar across multiple time steps within an episode [^1] While profiling the update ratios, we observed substantial variability in the visible detection module. This counterintuitive result stems from the instability inherent in the visual reasoning module of ECoT and the open-vocabulary nature of object labels, whereby identical objects may be assigned varying labels over time.. For example, the planning module exhibits an average update ratio of only $8.4\%$, meaning $91.6\%$ of its reasoning content is unchanged and can be reused in subsequent inference steps without autoregressive regeneration. Leveraging this temporal locality, we propose to cache ECoT reasoning for reuse in successive time steps, thus potentially enabling the parallel generation of each reasoning step.
@@
Figure 3: Comparison between ECoT (left) and our proposed Fast ECoT (right). Both decompose reasoning into fixed stages (e.g., task, plan, object grounding), but ECoT generates these sequentially at every step, while Fast ECoT enables parallel generation and reuses cached high-level reasoning from previous timesteps. The dotted lines coloured in green/magenta represent token copying.
4.2 Efficient Parallelisation of ECoT Reasoning and Action Generation
@@
Compared to the original ECoT—which generates the full reasoning sequence sequentially and autoregressively at every timestep (see Fig. 3 left)—our Fast ECoT reformulates each reasoning step $r_{n}^{t}$ as a standalone generation task. For each step, we construct the input by combining the current observation $O^{t}$, instruction $I^{t}$, and the previously generated reasoning steps $R^{t}=\{r_{i}^{t}\mid i=1,2,\dots,n-1\}$ from the last timestep as prefix context. This allows all reasoning steps and the action at timestep $t$ to be generated independently and in parallel, rather than waiting for preceding components to finish (see Fig. 3 right).
@@
To address inefficiencies from padding, we adopt continuous batching [22], a dynamic scheduling strategy used in modern LLM serving engines[24, 23]. Instead of padding all sequences to a fixed length, continuous batching maintains a queue where completed sequences are immediately replaced by new ones, allowing variable-length inputs to be processed efficiently. This minimises wasted computation on padding tokens, significantly improves GPU utilisation, and can reduce the total number of tokens processed by up to 6× (see Fig. 4). We use vLLM [24] as our inference backend, and the full pseudocode for Fast ECoT is provided in Sec. A.2.
Figure 4: Illustrative example comparing static vs. continuous batching for reasoning generation. Left: Static batching pads to the longest sequence, processing 4×11=44 tokens. Right: Continuous batching processes only actual tokens (3, 6, 8, 9), adding up to 26 tokens, reducing padding and improving efficiency.
4.3 Optimising Asynchronous Scheduling of Reasoning and Action Updates
@@
As discussed in Sec. 4.2, we formulate reasoning and action generation at each time step as batched requests that can be processed in parallel. However, reasoning traces typically span hundreds of tokens, while action decoding involves only a few (around $7$) tokens. In a synchronised setup, this mismatch causes unnecessary latency: the agent must wait for all reasoning steps to complete before it can act.
While the original ECoT framework couples reasoning and action tightly for interpretability and causal grounding, we argue that this coupling need not be time-synchronous. In practice, high-level reasoning elements—such as Task and Plan—remain stable across multiple steps, and their semantic influence on action decisions persists even when updated less frequently. This is evident in the rollout shown in Fig. 5, where reasoning traces and visible objects stay largely consistent despite continuously evolving actions. Motivated by this temporal stability, we adopt an asynchronous design that reuses cached high-level reasoning across timesteps, avoiding unnecessary regeneration. This is particularly beneficial in robotic settings, where action frequencies are faster than reasoning updates. By decoupling reasoning from action generation and updating the cached reasoning trace $R^{c}$ in the background, our approach enables low-latency action decoding without compromising interpretability or task performance. The asynchronous scheduling mechanism separately handles reasoning and action requests, returning actions immediately upon decoding, while reasoning is refreshed opportunistically. Full pseudocode is provided in Alg. 1.
@@ 5 Experimental Results
5.1 LIBERO Simulation Experimental Setup
@@
Training Data and Training Recipe. ECoT models require fine-tuning on each task for optimal performance [1, 4]. Training data is generated following the ECoT pipeline [4], with modifications for reproduction: we integrate Janus [30] for automated scene description and Deepseek-Reasoner [26] for generating CoT reasoning trajectories from successful demonstrations. We initialise ECoT models from the open-sourced checkpoint [4], which was pre-trained on Bridge V2 [5] and OXE dataset [6]. Then we apply LoRA [31] with rank 32 and train for $200,000$ gradient steps using a batch size of 1 distributed across 4 NVIDIA A6000 GPUs. See Sec. B.1 for other hyperparameters.
@@ 5.2 LIBERO Simulation Experimental Results
ECoT 77 84 75 45 70.3 4997 $\pm$ 691
ECoT (5-step) 79 75 72 40 66.5 3514 $\pm$ 969
ECoT (Async) 70 83 80 51 71.0 3655 $\pm$ 773
ECoT (Quant) 82 82 84 53 75.3 2180 $\pm$ 207
@@
Table 1: LIBERO simulation experimental results. SR = Success Rate. All results are averaged over 100 trials per suite and run on a single NVIDIA RTX 4090 GPU. Best results are in bold.
Inference Speedup. As shown in the last column of Tab. 1, our parallel reasoning strategy significantly accelerates inference latency. Specifically, Fast ECoT achieves a latency reduction to $2156\pm 353$ ms per step, representing over a 2.3x speed-up compared to the original ECoT baseline ($4997\pm 691$ ms). The asynchronous variant, Fast ECoT (Async), further reduces latency to $686\pm 412$ ms, nearly a 7x improvement. Notably, both variants maintain relatively stable latency compared to baselines proposed in ECoT [4].
Task Success Rate. Despite the significant reduction in latency, our method preserves or enhances task performance relative to baseline methods (see Tab. 1). Fast ECoT achieves the highest overall performance, with an average success rate (SR) of $80.0\%$, exceeding the baseline ECoT in all LIBERO suites. We hypothesise that our parallel reasoning strategy—which incorporates prior reasoning steps—helps smooth temporal inconsistencies and contributes to this performance gain. Notably, Fast ECoT (Async) also achieves a comparable overall performance, with an average SR of $77.5\%$, including the best result on LIBERO-Long ($69\%$). The largest decline compared with normal Fast ECoT happens at LIBERO-Object (from 83% to 75%), where the testing scenes focus on diverse layouts of the object. The positions of visible objects lead to a greater impact on overall control performance, and the larger temporal mismatch between reasoning components and actions of Fast ECoT (Async) may cause this degradation. Surprisingly, ECoT (Quant) also shows decent performance improvements compared to the original ECoT, which can be understood as regularising model behaviour in testing environments [34].
5.3 Real-world Experiments
Setup. We validate Fast ECoT on a physical Franka Emika Panda robotic arm equipped with a RealSense D455 camera, providing third-person RGB-D observations. We design six manipulation tasks representative of common household scenarios. The evaluation includes both in-distribution and out-of-distribution tasks featuring unseen objects and instructions, detailed in Sec. B.2. We collect 50 expert demonstrations via Droid teleoperation pipeline [35] to fine-tune ECoT models. We follow the same training data generation process, training recipe, and baselines used in Sec. 5.1.

### v2

@@ Abstract
Embodied Chain-of-Thought (ECoT) reasoning enhances vision-language-action (VLA) models by improving performance and interpretability through intermediate reasoning steps. However, its sequential autoregressive token generation introduces significant inference latency, limiting real-time deployment. We propose Fast ECoT, an inference-time acceleration method that exploits the structured and repetitive nature of ECoT to (1) cache and reuse high-level reasoning across timesteps and (2) parallelise the generation of modular reasoning steps. Additionally, we introduce an asynchronous scheduler that decouples reasoning from action decoding, further boosting responsiveness. Fast ECoT requires no model changes or additional training and easily integrates into existing VLA pipelines. Experiments in both simulation (LIBERO) and real-world robot tasks show up to a 7.5× reduction in latency with comparable or improved task success rate and reasoning faithfulness, bringing ECoT policies closer to practical real-time deployment. The code will be released upon acceptance.
I INTRODUCTION
@@
Building on this reuse, we introduce a partial parallelisation strategy that transforms the traditionally sequential reasoning process into a batched one. By branching only when necessary to handle novel information, Fast ECoT enables multiple reasoning steps to be generated in parallel—substantially reducing inference latency. We further introduce an asynchronous scheduling mechanism that decouples action and reasoning generation. Recognising that robot actions evolve faster than reasoning traces, we prioritise fast action decoding while allowing reasoning traces to update asynchronously in the background. This design reduces latency without compromising decision quality, as the cached reasoning remains stable over short time horizons.
II RELATED WORK
@@
Foundation Models for Robotic Manipulation. Recent advances in robot learning have produced large-scale generalist policies that excel at robotic manipulation tasks, by first pre-training on diverse multimodal datasets [1, 2, 3], and further fine-tuning on extensive robot-specific data collections [5, 6]. Vision-language-action (VLA) architectures [1, 2, 3]—which integrate vision-language models pretrained on internet-scale corpora [7, 8]—unify perception, language, and control into a single transformer-based policy, achieving state-of-the-art performance.
Current VLA methods can be broadly categorized into two families: (1) monolithic models, which integrate perception, language, and action within either single- or dual-system architectures [1, 2, 3]; and (2) hierarchical models, which explicitly decouple planning from policy execution by producing interpretable intermediate representations such as subtasks, keypoints, or programs [9, 10]. While monolithic approaches highlight simplicity and end-to-end generalisation, hierarchical designs emphasise interpretability and modularity for long-horizon control.
Reasoning for Robotic Control. Chain-of-thought (CoT) prompting [11, 12] has proven effective in enhancing large language models by encouraging step-by-step reasoning. Prior methods employ pre-trained language models for high-level planning [13, 14, 15], often requiring separate low-level controllers for execution. Recent work begins to integrate explicit reasoning end-to-end: ECoT introduces embodied chain-of-thought traces grounded in observations [4], CoT-VLA predicts visual subgoal observations [16], and RAD [17] and ThinkAct [18] curate or align language reasoning for low-level action; however, these often rely on latent embeddings, generated sub-goals, or textual descriptions that are hard to precisely ground for manipulation. EMMAX [19] embeds reasoning as subtasks and predicted gripper states (2D/3D), but leverages limited full-scene context. In contrast, MolmoAct [20] performs “reasoning in space,” explicitly grounding each step in the scene so it can be decoded and visualised on the image plane and within the 3D environment, improving explainability and action prediction. Orthogonal to reasoning design, efficiency-oriented VLAs such as Spec-VLA [21] and FlashVLA [22] boost responsiveness by replacing strictly sequential decoding with speculative, parallel, or retraining-free acceleration. Our work builds on the ECoT but targets its chief bottleneck—autoregressive latency—via a partially parallelised reasoning framework that preserves interpretability while substantially reducing inference time.
Inference Optimisation in Language and Multimodal Models. A wide range of techniques has been proposed to speed up inference in autoregressive models. Speculative decoding [23, 24] accelerates generation by predicting tokens with a lightweight draft model and verifying them with the full model. Non-autoregressive and parallel decoding strategies have also been adopted, particularly in machine translation and, more recently, in robotic control [9, 25]. Quantisation methods [26, 27, 28] are widely used to reduce model precision for faster computation. In contrast to these methods, which typically operate at the token or model level, our work focuses on reasoning-level acceleration—breaking CoT generation into reusable, semantically meaningful segments that can be cached and executed in parallel. This approach is model-agnostic and does not rely on auxiliary models, fine-tuning, or precision reduction. By enabling efficient reuse and branching during reasoning, our method offers a lightweight and integrable solution for speeding up VLA policies with CoT reasoning, without compromising interpretability or task success.
III PRELIMINARIES
III-A Embodied Chain-of-Thought Reasoning (ECoT)
@@
Vision‑Language‑Action (VLA) models [1] build on large pre‑trained vision‑language models, fine‑tuning them to map a natural language instruction $I^{t}$ and image observation $O^{t}$ directly to low‑level robot actions $A^{t}$ via autoregressive token prediction at each control step $t$. Embodied Chain‑of‑Thought (ECoT) [4] reasoning augments this reactive paradigm by supervising the model to generate a structured sequence of $N$ intermediate reasoning steps $R^{t}=\{\,r_{i}^{t}\mid i=1,2,\dots,N\}$, e.g. rephrased task, high‑level plan, grounded sub‑task, low‑level move command, visual features, before emitting the final action $A^{t}$. The reasoning steps are separated into high- and low-level steps (see Fig. 1).
III-B Continuous Batching
@@
In autoregressive models, batching improves throughput, but static batching can be inefficient when sequence lengths vary—shorter sequences must wait for longer ones, wasting compute on padding. Continuous batching [29] addresses this by dynamically replacing completed sequences with new ones, maintaining high GPU utilization and minimising idle time. This strategy, adopted by the work [30, 31], has shown 2–4× throughput gains in LLM serving.
IV Method
(a) Mean update ratio ($\pm$1 standard deviation) per reasoning step, showing high-level stability versus frequent low-level updates.
@@ IV Method
(b) Mean length ($\pm$1 standard deviation) of different reasoning steps, showing significant disparities in token counts.
@@
Figure 2: Statistics illustrating the pattern of ECoT reasoning steps under Libero Goal [32] and Bridge V2 [5].
IV-A Inference Characteristics of ECoT Reasoning
@@
Unlike traditional chain-of-thought approaches in language modelling [33, 34] and vision-language modelling [35, 36], which yield diverse, dynamic reasoning patterns that rarely recur, ECoT consistently follows a structured, periodically recurring workflow: planning, sub-task identification, motion reasoning, and visual-feature processing, before predicting subsequent robot actions. To analyse this behaviour quantitatively, we sample all episodes containing reasoning from the Bridge V2 dataset [5] and compute the average and the standard deviation of the update ratio (percentage of reasoning content updated at the next time step) and token length for each reasoning step. As depicted in Fig. 2(a), higher-level reasoning components in ECoT (e.g., planning and subtask reasoning) remain relatively similar across multiple time steps within an episode [^1] While profiling the update ratios, we observed substantial variability in the visible detection module. This counterintuitive result stems from the instability inherent in the visual reasoning module of ECoT and the open-vocabulary nature of object labels, whereby identical objects may be assigned varying labels over time.. For example, the planning module exhibits an average update ratio of only $8.4\%$, meaning $91.6\%$ of its reasoning content is unchanged and can be reused in subsequent inference steps without autoregressive regeneration. Leveraging this temporal locality, we propose to cache ECoT reasoning for reuse in successive time steps, thus potentially enabling the parallel generation of each reasoning step. We observe a similar trend in the simulated environment, LIBERO-Goal [32]: high-level reasoning updates rarely while low-level components refresh more frequently (see Fig. 2); This consistency supports caching high-level reasoning and densely updating low-level content.
@@
Figure 3: Comparison between ECoT (left) and our proposed Fast ECoT (right). Both decompose reasoning into fixed stages (e.g., task, plan, object grounding), but ECoT generates these sequentially at every step, while Fast ECoT enables parallel generation and reuses cached higher-level reasoning from previous timesteps as context. The dotted lines coloured in green/magenta represent token copying.
IV-B Parallelising Reasoning and Action Generation
@@
Compared to the original ECoT—which generates the full reasoning sequence sequentially and autoregressively at every timestep (see Fig. 3 left)—our Fast ECoT reformulates each reasoning step $r_{n}^{t}$ as a standalone generation task. For each step, we construct the input by combining the current observation $O^{t}$, instruction $I^{t}$, and the previously generated reasoning steps $R^{t-1}=\{r_{i}^{t-1}\mid i=1,2,\dots,n-1\}$ from the last timestep as prefix context. This allows all reasoning steps and the action at timestep $t$ to be generated independently and in parallel, rather than waiting for preceding components to finish (see Fig. 3 right).
@@
To address inefficiencies from padding, we adopt continuous batching [29], a dynamic scheduling strategy used in modern LLM serving engines [30, 31]. Instead of padding all sequences to a fixed length, continuous batching maintains a queue where completed sequences are immediately replaced by new ones, allowing variable-length inputs to be processed efficiently. This minimises wasted computation on padding tokens, significantly improves GPU utilisation, and can reduce the total number of tokens processed by up to 6× (see Fig. 4). We use vLLM [31] as our inference backend, and the pseudocode of Fast ECoT is shown in Alg. 1.
Algorithm 1 Parallel Embodied Chain‑of‑Thought
[…]
4:for $i=1\,\mathbf{to}\,N+1$ concurrently do
6:end for
7:Synchronize
[…]
9:return $R^{t},A^{t}$
Figure 4: Illustratively comparing static vs. continuous batching for reasoning generation. Left: Static batching pads to the longest sequence, processing 4×11=44 tokens. Right: Continuous batching processes only actual tokens (3, 6, 8, 9), adding up to 26 tokens, which reduces padding and improves efficiency.
IV-C Asynchronous Reasoning and Action Updates
So far we formulate reasoning and action generation at each time step as batched requests that can be processed in parallel. However, reasoning traces typically span hundreds of tokens, while action decoding involves only a few (around $7$) tokens. In a synchronised setup, this mismatch causes unnecessary latency: the agent must wait for all reasoning steps to complete before it can act (see Fig. 6).
@@ IV-C Asynchronous Reasoning and Action Updates
(a) Synchronous inference.
(b) Asynchronous inference.
Figure 6: Illustration of inference stacks. Asynchronous overlap reduces stall time (green dashed) of robot operation and increases action throughput within the same time window compared to synchronous execution.
While ECoT couples reasoning and action for interpretability and causal grounding, this coupling need not be time-synchronous. Empirically, high-level elements (e.g., Task, Plan) change slowly across steps; their influence on actions persists even with infrequent updates—see rollouts in Fig. 5, where reasoning traces and visible objects remain stable as actions evolve. Inspired by VLA systems that separate a fast action module from a slower VLM (SmolVLA [37]; GR00T [38]), we adopt an asynchronous split: the controller decodes actions from the current observation $O^{t}$ and a cached high-level reasoning $R^{c}$, while $R^{c}$ is refreshed in the background. As shown in Fig. 6, asynchronous reasoning action generation effectively reduces stalls and increases action throughput without sacrificing interpretability. Full pseudocode of the above is shown in Alg. 2.
@@ V RESULTS
V-A LIBERO Experiments
@@
Training Data and Training Recipe. ECoT models require fine-tuning on each task for optimal performance [1, 4]. Training data is generated following the ECoT pipeline [4], with modifications for reproduction: we integrate Janus [39] for automated scene description and Deepseek-Reasoner [34] for generating CoT reasoning trajectories from successful demonstrations. We initialise ECoT models from the open-sourced checkpoint [4], which was pre-trained on Bridge V2 [5] and OXE dataset [6]. Then we apply LoRA [40] with rank 32 and train for $200,000$ gradient steps using a batch size of 1 distributed across 4 NVIDIA A6000 GPUs.
@@
OpenVLA 87 83 74 55 75.3 184 $\pm$ 37
ECoT 77 84 75 57 73.3 4997 $\pm$ 691
ECoT (5-step) 79 75 72 56 70.5 3514 $\pm$ 969
ECoT (Async) 70 83 80 47 70.0 3655 $\pm$ 773
ECoT (Quant) 82 82 84 57 76.3 2180 $\pm$ 207
@@
TABLE I: LIBERO simulation experimental results. SR = Success Rate.
[…]
Performance. Our method consistently improves task performance relative to prior baselines. Fast ECoT achieves the highest average success rate ($80.0\%$), surpassing all ECoT variants and non-reasoning OpenVLA. We attribute this to our parallel reasoning strategy, which smooths temporal inconsistencies and leverages prior reasoning steps. Fast ECoT (Async) remains competitive ($77.5\%$) and achieves the best result on LIBERO-Long ($69\%$), though at some cost in LIBERO-Object due to greater sensitivity to object layouts and temporal mismatch. Interestingly, ECoT (Quant) also improves over vanilla ECoT, likely due to quantisation regularising model behaviour [43]. Fig. 7 shows that both ECoT and Fast ECoT (Async) produce coherent, grounded reasoning, but the latter reaches comparable states much quicker.
(a) ECoT at $t=143.7$s.
(b) Fast ECoT (Async) at $t=21.4$s.
Figure 7: Qualitative examples in LIBERO experiments. Fast ECoT (Async) reaches comparable task states much quicker.
V-B Real-world Experiments
Setup. We validate Fast ECoT on a physical Franka Emika Panda robotic arm equipped with a RealSense D455 camera, providing third-person RGB-D observations. We design six manipulation tasks representative of common household scenarios. The evaluation includes both in-distribution and out-of-distribution tasks featuring unseen objects and instructions.We collect 50 expert demonstrations via Droid teleoperation pipeline [44] to fine-tune ECoT models. We follow the same training data generation process, training recipe, and baselines used in Sec. V-A.